# Three-dimensional fracton topological orders with boundary Toeplitz braiding

### v4

@@ A.2 $\det B\neq 0$
The boundary bases corresponding to $\beta_{2}$ and $\beta_{3}$ are denoted as $\mathbf{w}^{2}$ and $\mathbf{w}^{3}$, respectively.
$\mathbf{w}^{2}=\begin{pmatrix}\beta_{2}w^{2,a}&\beta_{2}w^{2,b}&\beta_{2}^{2}w^{2,a}&\beta_{2}^{2}w^{2,b}&\cdots&\beta_{2}^{N}w^{2,a}&\beta_{2}^{N}w^{2,b}\end{pmatrix}^{T},$
$\mathbf{w}^{3}=\begin{pmatrix}\beta_{3}w^{3,a}&\beta_{3}w^{3,b}&\beta_{3}^{2}w^{3,a}&\beta_{3}^{2}w^{3,b}&\cdots&\beta_{3}^{N}w^{3,a}&\beta_{3}^{N}w^{3,b}\end{pmatrix}^{T}.$
The bulk recurrence relation Eq. (86) gives
The solution to Eq. (157) is given by
A trial solution $\mathbf{u}^{2,3}$ is a linear combination of $\mathbf{w}^{2}$ and $\mathbf{w}^{3}$.
where
Since $|\beta_{2}|<1,|\beta_{3}|<1$, $\mathbf{u}^{2,3}$ is required to satisfy the boundary condition specified by Eq. (94). If $\beta_{2}\neq\beta_{3}\Leftrightarrow l_{2}\neq l_{3}$, the boundary condition Eq. (94) renders
Nontrivial solution to Eq. (160) implies
Eq. (163) together with Eq. (146b), Eq. (146c) renders
$2ml_{2}\beta_{2}+2l_{2}(l_{2}+l_{3})=2ml_{3}\beta_{3}+2l_{2}(l_{2}+l_{3})=0,$ (164)
Eq. (164) together with Eq. (146b) and Eq. (146c) enforces $\beta_{2}=1$, $\beta_{3}=l_{2}/l_{3}$, which contradicts our assumption $|\beta_{2}|<1,|\beta_{3}|<1$. If $\beta_{2}=\beta_{3}\Leftrightarrow l_{2}=l_{3}$, a legitimate solution $\mathbf{u}^{2,3}$ satisfying the boundary condition Eq. (94) does not exist either.
The next step is to discuss the boundary bases corresponding to $\beta_{1}$ and $\beta_{4}$, which are denoted as $\mathbf{w}^{1}$ and $\mathbf{w}^{4}$, respectively.
$\mathbf{w}^{1}=\begin{pmatrix}\beta_{1}w^{1,a}&\beta_{1}w^{1,b}&\beta_{1}^{2}w^{1,a}&\beta_{1}^{2}w^{1,b}&\cdots&\beta_{1}^{N}w^{1,a}&\beta_{1}^{N}w^{1,b}\end{pmatrix}^{T},$
$\mathbf{w}^{4}=\begin{pmatrix}\beta_{4}w^{4,a}&\beta_{4}w^{4,b}&\beta_{4}^{2}w^{4,a}&\beta_{4}^{2}w^{4,b}&\cdots&\beta_{4}^{N}w^{4,a}&\beta_{4}^{N}w^{4,b}\end{pmatrix}^{T}.$
@@
A trial solution $\mathbf{u}^{1,4}$ is a linear combination of $\mathbf{w}^{1}$ and $\mathbf{w}^{4}$.
@@
Since $|\beta_{1}|>1,|\beta_{4}|>1$, $\mathbf{u}^{1,4}$ is required to satisfy the boundary condition Eq. (94). If $\beta_{2}\neq\beta_{3}\Leftrightarrow l_{2}\neq l_{3}$, the boundary condition Eq. (95) renders
@@
Eq. (172) together with Eq. (146a) and Eq. (146d) enforces $\beta_{4}=1$, $\beta_{1}=l_{2}/l_{3}$, which contradicts our assumption $|\beta_{1}|>1,|\beta_{4}|>1$. If $\beta_{1}=\beta_{4}\Leftrightarrow l_{2}=l_{3}$, a legitimate solution $\mathbf{u}^{1,4}$ satisfying the boundary condition Eq. (95) does not exist either.
@@
Denote the boundary bases corresponding to $\beta_{1}$ and $\beta_{4}$ as $\mathbf{w}^{1}$ and $\mathbf{w}^{4}$, respectively.
$\mathbf{w}^{1}=\begin{pmatrix}\beta_{1}w^{1,a}&\beta_{1}w^{1,b}&\beta_{1}^{2}w^{1,a}&\beta_{1}^{2}w^{1,b}&\cdots&\beta_{1}^{N}w^{1,a}&\beta_{1}^{N}w^{1,b}\end{pmatrix}^{T},$
$\mathbf{w}^{4}=\begin{pmatrix}\beta_{4}w^{4,a}&\beta_{4}w^{4,b}&\beta_{4}^{2}w^{4,a}&\beta_{4}^{2}w^{4,b}&\cdots&\beta_{4}^{N}w^{4,a}&\beta_{4}^{N}w^{4,b}\end{pmatrix}^{T}.$
The bulk recurrence relation Eq. (86) gives
Eq. (173) is solved by
A trial solution is a linear combination of $\mathbf{w}^{1}$ and $\mathbf{w}^{4}$.
where
Since $|\beta_{1}|<1,|\beta_{4}|<1$, $\mathbf{u}^{1,4}$ is required to satisfy the boundary condition specified by Eq. (94). If $\beta_{1}\neq\beta_{4}\Leftrightarrow l_{2}\neq l_{3}$, the boundary condition Eq. (94) renders
Nontrivial solution to Eq. (176) requires
Eq. (179) along with Eq. (146a) and Eq. (146d) renders
Eq. (180), Eq. (146a) and Eq. (146d) enforce $\beta_{4}=1$, $\beta_{1}=l_{3}/l_{2}$, which contradicts our assumption $|\beta_{1}|<1$, $|\beta_{4}|<1$. If $\beta_{1}=\beta_{4}\Leftrightarrow l_{2}=l_{3}$, a legitimate solution $\mathbf{u}^{1,4}$ satisfying the boundary condition Eq. (94) does not exist either.
The next step is to discuss the boundary bases $\mathbf{w}^{2}$ and $\mathbf{w}^{3}$ corresponding to $\beta_{2}$ and $\beta_{3}$ as, respectively.
$\mathbf{w}^{2}=\begin{pmatrix}\beta_{2}w^{2,a}&\beta_{2}w^{2,b}&\beta_{2}^{2}w^{2,a}&\beta_{2}^{2}w^{2,b}&\cdots&\beta_{2}^{N}w^{2,a}&\beta_{2}^{N}w^{2,b}\end{pmatrix}^{T},$
$\mathbf{w}^{3}=\begin{pmatrix}\beta_{3}w^{3,a}&\beta_{3}w^{3,b}&\beta_{3}^{2}w^{3,a}&\beta_{3}^{2}w^{3,b}&\cdots&\beta_{3}^{N}w^{3,a}&\beta_{3}^{N}w^{3,b}\end{pmatrix}^{T}.$
The bulk recurrence relation Eq. (86) gives
@@
A trial solution is a linear combination of $\mathbf{w}^{2}$ and $\mathbf{w}^{3}$.
@@
Since $|\beta_{2}|>1,|\beta_{3}|>1$, $\mathbf{u}^{2,3}$ is required to satisfy the boundary condition specified by Eq. (95). If $\beta_{2}\neq\beta_{3}\Leftrightarrow l_{2}\neq l_{3}$, the boundary condition (95) renders
@@
Eq. (189), Eq. (146b) and Eq. (146c) enforce $\beta_{2}=1$, $\beta_{3}=l_{3}/l_{2}$, which contradicts our assumption $|\beta_{2}|>1$, $|\beta_{3}|>1$. If $\beta_{2}=\beta_{3}\Leftrightarrow l_{2}=l_{3}$, it is easy to verify that a legitimate boundary zero mode satisfying the boundary condition Eq. (95) does not exist either.
$|\beta_{1}|>1$, $|\beta_{2}|<1$, $|\beta_{3}|>1$, $|\beta_{4}|<1$ and $|\beta_{1}|<1$, $|\beta_{2}|>1$, $|\beta_{3}|<1$, $|\beta_{4}|>1$ can be summarized into the condition $|m|<|l_{2}+l_{3}|$. If $\left\{\begin{array}[]{l}|m|<|l_{2}+l_{3}|\\ |l_{3}|<|l_{2}|\end{array}\right.$, the discussion falls in Case I. If $\left\{\begin{array}[]{l}|m|<|l_{2}+l_{3}|\\ |l_{2}|<|l_{3}|\end{array}\right.$, the discussion falls in Case II.
In summary, if
$\left\{\begin{array}[]{l}l_{4}=0,\\ n(l_{2}+l_{3})\neq 2l_{1}m,\quad l_{2}\neq l_{3},\quad l_{2}l_{3}\neq 0,\\ |m|<|l_{2}+l_{3}|,\end{array}\right.$
the $K$-matrix Eq. (2) has one boundary zero mode. More specifically, if $|l_{2}|<|l_{3}|$, the boundary zero mode is a lower boundary zero mode, satisfying the boundary condition Eq. (94). If $|l_{2}|>|l_{3}|$, the boundary zero mode is an upper boundary zero mode, satisfying the boundary condition Eq. (95). If
$\left\{\begin{array}[]{l}l_{4}=0,\quad n(l_{2}+l_{3})=2l_{1}m,\\ l_{2}\neq l_{3},\quad l_{2}l_{3}\neq 0,\\ |m|<|l_{2}+l_{3}|,\end{array}\right.$
the $K$-matrix Eq. (2) has two boundary zero modes.

### v5

@@ Appendix C The Conditions for the Existence of Boundary Zero Modes under $\det B\neq 0$
Since $|\beta_{1}|>1,|\beta_{4}|>1$, $\mathbf{u}^{1,4}$ is required to satisfy the boundary condition Eq. (151a). The boundary bases corresponding to $\beta_{1}$ and $\beta_{4}$ are denoted as $\mathbf{w}^{1}$ and $\mathbf{w}^{4}$, respectively.
$\mathbf{w}^{j}=\begin{pmatrix}\beta_{j}w^{j,a}&\beta_{j}w^{j,b}&\beta_{j}^{2}w^{j,a}&\beta_{j}^{2}w^{j,b}&\cdots&\beta_{j}^{N}w^{j,a}&\beta_{j}^{N}w^{j,b}\end{pmatrix}^{T},\quad j=1,4.$
@@
$\mathbf{u}^{1,4}$ is a linear combination of $\mathbf{w}^{1}$ and $\mathbf{w}^{4}$.
@@
The boundary condition Eq. (151b) renders
@@
Eq. (360) together with Eq. (305a) and Eq. (305d) enforces $|\beta_{4}|=1$, which contradicts our assumption $|\beta_{1}|>1,|\beta_{4}|>1$.
@@
Since $|\beta_{2}|>1,|\beta_{3}|>1$, $\mathbf{u}^{2,3}$ is required to satisfy the boundary condition specified by Eq. (151b). The boundary bases corresponding to $\beta_{2}$ and $\beta_{3}$ are denoted as $\mathbf{w}^{2}$ and $\mathbf{w}^{3}$, respectively.
$\mathbf{w}^{j}=\begin{pmatrix}\beta_{j}w^{j,a}&\beta_{j}w^{j,b}&\beta_{j}^{2}w^{j,a}&\beta_{j}^{2}w^{j,b}&\cdots&\beta_{j}^{N}w^{j,a}&\beta_{j}^{N}w^{j,b}\end{pmatrix}^{T},\quad j=2,3.$
The bulk recurrence relation Eq. (144) yields
@@
$\mathbf{u}^{2,3}$ is a linear combination of $\mathbf{w}^{2}$ and $\mathbf{w}^{3}$.
@@
The boundary condition (151b) renders
@@
Eq. (369), Eq. (305b) and Eq. (305c) enforce $|\beta_{2}|=1$, which contradicts our assumption $|\beta_{2}|>1$, $|\beta_{3}|>1$.
C.3 The conditions for boundary zero modes under $\det B\neq 0,l_{4}=0$
The set operation $\mathbb{C}_{1}\cup\mathbb{C}_{2}$ yields the condition for the appearance of a boundary zero mode satisfying the boundary condition Eq. (151a) under $\det B\neq 0$ and $l_{4}=0$:
$\mathbb{C}_{1}:\ \left\{\begin{array}[]{l}l_{4}=0,\quad n(l_{2}+l_{3})=2l_{1}m,\\ m\neq 0,\quad l_{2}l_{3}\neq 0,\\ |m|<|l_{2}+l_{3}|,\quad|l_{3}|<|l_{2}|,\\ \end{array}\right.\quad\text{or}\quad\mathbb{C}_{2}:\ \left\{\begin{array}[]{l}l_{4}=0,\\ m\neq 0,\quad l_{2}l_{3}\neq 0,\\ |m|<|l_{2}+l_{3}|,\quad|l_{2}|<|l_{3}|.\end{array}\right.$ (376)
Here, the original condition $|\beta_{1}|>1$, $|\beta_{2}|<1$, $|\beta_{3}|>1$, $|\beta_{4}|<1$ simplifies to
$\left\{\begin{array}[]{l}|m|<|l_{2}+l_{3}|,\\ |l_{3}|<|l_{2}|,\end{array}\right.$
and the original condition $|\beta_{1}|<1$, $|\beta_{2}|>1$, $|\beta_{3}|<1$, $|\beta_{4}|>1$ simplifies to
$\left\{\begin{array}[]{l}|m|<|l_{2}+l_{3}|,\\ |l_{2}|<|l_{3}|.\end{array}\right.$
Likewise, the condition for the appearance of a boundary zero mode satisfying the boundary condition Eq. (151b) given by the set operation $\mathbb{D}_{1}\cup\mathbb{D}_{2}$ reads
$\mathbb{D}_{1}:\ \left\{\begin{array}[]{l}l_{4}=0,\\ m\neq 0,\quad l_{2}l_{3}\neq 0,\\ |m|<|l_{2}+l_{3}|,\quad|l_{3}|<|l_{2}|,\end{array}\right.\quad\text{or}\quad\mathbb{D}_{2}:\ \left\{\begin{array}[]{l}l_{4}=0,\quad n(l_{2}+l_{3})=2l_{1}m,\\ m\neq 0,\quad l_{2}l_{3}\neq 0,\\ |m|<|l_{2}+l_{3}|,\quad|l_{2}|<|l_{3}|.\\ \end{array}\right.$ (383)
The condition for the existence of boundary zero modes is denoted as $\mathbb{U}=\mathbb{C}_{1}\cup\mathbb{C}_{2}\cup\mathbb{D}_{1}\cup\mathbb{D}_{2}$. The condition for the appearance of exactly one boundary zero mode is
$\left\{\begin{array}[]{l}l_{4}=0,\\ n(l_{2}+l_{3})\neq 2l_{1}m,\quad l_{2}\neq l_{3},\quad l_{2}l_{3}\neq 0,\\ |m|<|l_{2}+l_{3}|,\end{array}\right.$ (387)
derived from the set operation $\mathbb{U}\backslash[(\mathbb{C}_{1}\cup\mathbb{C}_{2})\cap(\mathbb{D}_{1}\cup\mathbb{D}_{2})]$. The zero mode manifests as either:
A lower-boundary zero mode satisfying Eq. (151a) when $|l_{2}|<|l_{3}|$, or
An upper-boundary zero mode satisfying Eq. (151b) when $|l_{2}|>|l_{3}|$.
The condition for two boundary zero modes is
$\left\{\begin{array}[]{l}l_{4}=0,\quad n(l_{2}+l_{3})=2l_{1}m,\\ l_{2}\neq l_{3},\quad l_{2}l_{3}\neq 0,\\ |m|<|l_{2}+l_{3}|,\end{array}\right.$ (391)
obtained from the complement of set $\mathbb{U}\backslash[(\mathbb{C}_{1}\cup\mathbb{C}_{2})\cap(\mathbb{D}_{1}\cup\mathbb{D}_{2})]$, i.e., $(\mathbb{C}_{1}\cup\mathbb{C}_{2})\cap(\mathbb{D}_{1}\cup\mathbb{D}_{2})$.